# Quantum Algorithms in a Superposition of Spacetimes

### v2

@@ Abstract
We show that a quantum computer with the ability to create a superposition of unitary orders is able to solve in polynomial time two fundamental problems in computer science: The Graph Isomorphism Problem ($\mathsf{GI}$) and the Gap Closest Vector Problem ($\mathsf{GapCVP}$), with gap $O\left(n^{2}\right)$. These problems are believed by experts to be hard to solve for a regular quantum computer. Interestingly, our model does not seem overpowered, and we found no obvious way to solve entire complexity classes that are considered hard in computer science, like the classes NP and SZK. As part of our work we develop a new parameterization technique for the invertibility of classical circuits, which raises new questions in the analysis of probability distributions and induces a new parameter for hardness inside the complexity class SZK. These new techniques are independent of any non-classical computation.
@@ 1.3 Results
We focus on two very well-studied computational tasks in computer science: The Graph Isomorphism problem ($\mathsf{GI}$, Definition 3.9) and the Gap Closest Vector Problem ($\mathsf{GapCVP}_{g(n)}$, Definition 3.11). Both $\mathsf{GI}$ and $\mathsf{GapCVP}_{g(n)}$ for $g(n):=O(n^{2})$ are believed by experts to be unsolvable in quantum polynomial time. In particular, $\mathsf{GapCVP}_{O(n^{2})}$ is a key problem in the theory of computation and have an immense practical significance: Solving $\mathsf{GapCVP}_{O(n^{2})}$ means solving the Learning With Errors ($\mathsf{LWE}$) problem [Reg09], the computational hardness of which is the central pillar of post-quantum cryptography. We prove the following theorems.
@@
There exists a positive absolute constant $c\in\mathbb{R}_{>0}$ such that $\mathsf{GapCVP}_{c\cdot n^{2}}\in\textbf{BQP}^{\text{OI}}$.
@@ 2 Technical Overview
In this section we give an overview of the main technical ideas in this work. In Section 2.1 we form an understanding of computational abilities that computable OI has and a standard quantum computer does not seem to posses. We also connect these understandings to the Statistical Difference problem. In Section 2.2 we show a new technique to parameterize invertibility of classical circuits and define a new computational problem, called the Sequentially Invertible Statistical Difference Problem $\mathsf{SISD}$. In Section 2.3 we show a polynomial-time quantum algorithm using an OI oracle, that solves $\mathsf{SISD}$. In Section 2.4 we show classical polynomial-time reductions from the Graph Isomorphism Problem $\mathsf{GI}$ and the $O(n^{2})$-Gap Closest Vector Problem $\mathsf{GapCVP}_{O\left(n^{2}\right)}$ to the $\mathsf{SISD}$ problem.
@@ 2.2 The Sequentially Invertible Statistical Difference Problem
Our definition of parameterized invertibility of a circuit is as follows. Let $C:\{0,1\}^{k^{\prime}}\rightarrow\{0,1\}^{k}$ a circuit. For $r\in\{0,1,\cdots,k^{\prime}\}$, we say that $D$ is $r$-sequentially invertible if there is an $r$-invertible circuit sequence that samples from the output distribution of $C$ [^7] The full definition of $\left(r,t,\ell,\varepsilon\right)$-sequential invertibility, which takes into consideration more parameters and not only $r$, is given in Section 4.3. For the needs for this overview, we give the minimal definitions.. An $r$-invertible circuit sequence on $k$ bits is a sequence of pairs of circuits $\left(C_{i,\rightarrow},C_{i,\leftarrow}\right)_{i\in[\ell]}$ such that for every $i\in[\ell]$,
@@
Similarly to the case of the regular Statistical Distance problem $\mathsf{SD}_{a(n),b(n)}$, we have the following promise on the gap of the statistical distance between the circuits. Formally, for $b\in\{0,1\}$ denote
@@
where the randomness sequence is sampled in both cases uniformly $z_{1}\leftarrow\{0,1\}^{r^{b}_{1}(n)},\cdots,z_{\ell}\leftarrow\{0,1\}^{r^{b}_{\ell}(n)}$ for the suitable $b\in\{0,1\}$. We denote the statistical distance between the distributions $D_{0}$, $D_{1}$ with $d$, and it is promised that either $d\leq a(n)$, in which case we say that the input belongs to $\prod_{\text{YES}}$, or we have $d>b(n)$, in which case we determine that the input belongs to $\prod_{\text{NO}}$. The goal of the problem is to decide whether the input belongs to $\prod_{\text{YES}}$ or to $\prod_{\text{NO}}$. As our previous convention for the $\mathsf{SD}$ problem, we also drop the subscript $a(n),b(n)$ to implicitly be any pair of functions with difference $b(n)^{2}-2a(n)+a(n)^{2}$ lower-bounded inverse polynomially, and write $\mathsf{SISD}_{\mathsf{poly},\,r(n)}$.
@@ 2.4 Reducing $\mathsf{GI}$ and $\mathsf{GapCVP}$ to $\mathsf{SISD}_{\mathsf{poly},O(\log(n))}$
In the last part of the Technical Overview we will show that two problems that are believed to be hard for quantum computers, are solvable by a quantum computer with computable order interference, by reducing to the $\mathsf{SISD}_{\mathsf{poly},O(\log(n))}$ problem. In the first part of this section we show the reduction from $\mathsf{GI}$ which is more immediate, and in the second part we show the more involved reduction from $\mathsf{GapCVP}_{O\left(n^{2}\right)}$. To this end, we will first review the existing reductions from each of the problems to the more general $\mathsf{SD}_{\mathsf{poly}}$ problem (which we don’t know how to solve), and then move to our new reductions, to the easier problem of $\mathsf{SISD}_{\mathsf{poly},O(\log(n))}$.
@@
  • $d\in\mathbb{R}_{\geq 0}$ is a non-negative real number.
@@
The input comes with a promise: It is either the case that the vector $\mathbf{t}$ is $d$-close to the lattice i.e., $\text{dist}\left(\mathcal{L}_{\mathbf{B}},\mathbf{t}\right)\leq d$, in which case we denote $\left(\mathbf{B},\mathbf{t},d\right)\in\prod_{\text{YES}}$, or the vector is farther than $d\cdot g(n)$ from the lattice i.e., $\text{dist}\left(\mathcal{L}_{\mathbf{B}},\mathbf{t}\right)>d\cdot g(n)$, in which case we denote $\left(\mathbf{B},\mathbf{t},d\right)\in\prod_{\text{NO}}$. In simple words, a solving algorithm for $\mathsf{GapCVP}_{g(n)}$ gets the promise that the vector $\mathbf{t}$ is either close or far away from the lattice $\mathcal{L}_{\mathbf{B}}$, and needs to decide which is the case. The bigger the gap $g(n)$ is, the easier the problem becomes [^9] In particular, it is known that for an exponential $g(n)\approx 2^{n}$ the $\mathsf{GapCVP}_{2^{n}}$ problem is solvable in classical polynomial time [AKS01].. What we will show is that for $g(n)=O\left(n^{2}\right)$ there is a reduction $\mathsf{GapCVP}_{g(n)}\leq_{p}\mathsf{SISD}_{\mathsf{poly},O(\log(n))}$. The inspiration and starting point of our reduction is the reduction $\mathsf{GapCVP}_{\sqrt{\frac{n}{O\left(\log\left(n\right)\right)}}}\leq_{p}\mathsf{SD}_{\mathsf{poly}}$ by Goldreich and Goldwasser [GG98], which we recall next.
@@
  • Density: We imagine the support of $P$ as an $n$-dimensional shape, by looking at the elements that are farthest away from the origin and thinking on them as a boundary for the shape. The distribution $P$ should be as dense as possible in its shape. The density requirement emerges from the case where $\left(\mathbf{B},\mathbf{t},d\right)\in\prod_{\text{YES}}$ – in that case we sample uniform samples from the shape around the lattice vectors, and the more the shape is sparse, the less our ability to give a good upper bound on the total variation distance between the output distributions of the circuits $C_{0}$, $C_{1}$.
@@
The distribution $P$ we choose is the $n$-dimensional discrete truncated Gaussian distribution $D^{n}_{B}$ (as per Definition 3.12, $B\in\mathbb{N}$ is the truncation parameter, assume it to be $\approx\frac{d\cdot g(n)}{2\cdot\sqrt{n}}$). It is a known fact that $D^{n}_{B}$ is invariant under rotations in $n$-space, which exactly means it does not prefer any direction of the $n$ directions in $\mathbb{Z}^{n}$. In terms of density, the Gaussian distribution is obviously less dense in its shape than the uniform $n$-ball, which is (part of) the reason that we need a bigger gap $g(n)=O\left(n^{2}\right)$ for which we know how to reduce $\mathsf{GapCVP}_{g(n)}$ to $\mathsf{SISD}$, compared to the better gap $\sqrt{\frac{n}{O\left(\log\left(n\right)\right)}}$ from the previous reduction. As part of the full proof of the reduction (given in Section 7) we analyze the Gaussian distribution’s sparsity in this manner, and the loss of hardness (which translates to the larger $g(n)\approx n^{2}>>\sqrt{\frac{n}{O\left(\log\left(n\right)\right)}}$) that the sparsity incurs.
@@
Sampling from $D_{B}$ has a plethora of existing techniques and algorithms (e.g., [GPV08, Pei10, MP12, Ros19]), but it is unclear how to show that these can be executed by a sequentially invertible algorithm. In fact, we did not find any sequentially invertible algorithm that samples from a distribution with negligible total variation distance to $D_{B}$. Once we relax our needs to sample from the discrete Gaussian distribution only approximately (that is, with a non-negligible total variation distance), we find a solution. We give the intuitive steps behind our solution in the following.
  • Recall one of the basic theorems of probability theory - The Central Limit Theorem (CLT). The CLT says that when we take the average of a collection of $\kappa$ i.i.d. random numerical variables $X_{1},\cdots,X_{\kappa}$, this average $\frac{\sum_{i\in[\kappa]}X_{i}}{\kappa}$ approaches the Gaussian distribution as $\kappa$ approaches infinity. This helps us because now, instead of sampling from the Gaussian distribution (which we do not know how to do with a sequentially invertible algorithm), we can repeatedly sample from any distribution. While the CLT asserts closeness to the Gaussian distribution for some number of summands, it does not guarantee a rate of convergence, as the rate depends on the chosen distribution of the summands $\{X_{i}\}_{i\in[\kappa]}$.
  • We observe that the uniform distribution has two properties that work in our favour, for the summands $\{X_{i}\}_{i\in[\kappa]}$. Using a strengthening of the CLT (a Berry-Esseen-type Theorem), it can be shown that the uniform distribution gives a good convergence rate to the Gaussian distribution, even in total variation distance and even in the discrete case. Formally, in Section 7 we show that summing $\kappa$ i.i.d. samples of the uniform distribution over $-2^{\beta},\cdots,-1,0,1,\cdots,2^{\beta}$ for $2^{\beta}\approx\frac{B}{\kappa}$ has total variation distance bounded by $O\left(\frac{1}{\sqrt{\kappa}}\right)$ to the discrete Gaussian distribution $D_{B}$.
@@
To conclude, the circuit sequence $C^{0}$ samples a uniformly random lattice vector $\mathbf{v}$, and then adds an approximate discrete Gaussian vector $\mathbf{e}$, where the approximation is done by summing i.i.d. samples from the uniform distribution. The circuit sequence $C^{1}$ is the same, only that we add $\mathbf{t}$ in the end. Both circuit sequences are shown to be $1$-sequentially invertible. We refer the reader to Section 7 for the full details of the proof, including the analysis of the total variation distance between the output distributions of the circuits.

### v3

@@ Abstract
We show that a quantum computer with the ability to create a superposition of unitary orders is able to solve in polynomial time two fundamental problems in computer science: The Graph Isomorphism Problem ($\mathsf{GI}$) and the Gap Closest Vector Problem ($\mathsf{GapCVP}$), with gap $O\left(n\sqrt{n}\right)$. These problems are believed by experts to be hard to solve for a regular quantum computer. Interestingly, our model does not seem overpowered, and we found no obvious way to solve entire complexity classes that are considered hard in computer science, like the classes NP and SZK. As part of our work we develop a new parameterization technique for the invertibility of classical circuits, which raises new questions in the analysis of probability distributions and induces a new parameter for hardness inside the complexity class SZK. These new techniques are independent of any non-classical computation.
@@ 1.3 Results
We focus on two very well-studied computational tasks in computer science: The Graph Isomorphism problem ($\mathsf{GI}$, Definition 3.9) and the Gap Closest Vector Problem ($\mathsf{GapCVP}_{g(n)}$, Definition 3.11). Both $\mathsf{GI}$ and $\mathsf{GapCVP}_{g(n)}$ for $g(n):=O\left(n\sqrt{n}\right)$ are believed by experts to be unsolvable in quantum polynomial time. In particular, $\mathsf{GapCVP}_{O\left(n\sqrt{n}\right)}$ is a key problem in the theory of computation and have an immense practical significance: Solving $\mathsf{GapCVP}_{O(n\sqrt{n})}$ means solving the Learning With Errors ($\mathsf{LWE}$) problem [Reg09], the computational hardness of which is the central pillar of post-quantum cryptography. We prove the following theorems.
@@
There exists a positive absolute constant $c\in\mathbb{R}_{>0}$ such that $\mathsf{GapCVP}_{c\cdot n\sqrt{n}}\in\textbf{BQP}^{\text{OI}}$.
@@ 2 Technical Overview
In this section we give an overview of the main technical ideas in this work. In Section 2.1 we form an understanding of computational abilities that computable OI has and a standard quantum computer does not seem to posses. We also connect these understandings to the Statistical Difference problem. In Section 2.2 we show a new technique to parameterize invertibility of classical circuits and define a new computational problem, called the Sequentially Invertible Statistical Difference Problem $\mathsf{SISD}$. In Section 2.3 we show a polynomial-time quantum algorithm using an OI oracle, that solves $\mathsf{SISD}$. In Section 2.4 we show classical polynomial-time reductions from the Graph Isomorphism Problem $\mathsf{GI}$ and the $O(n\sqrt{n})$-Gap Closest Vector Problem $\mathsf{GapCVP}_{O\left(n\sqrt{n}\right)}$ to the $\mathsf{SISD}$ problem.
@@ 2.2 The Sequentially Invertible Statistical Difference Problem
Our definition of parameterized invertibility of a circuit is as follows. Let $C:\{0,1\}^{k^{\prime}}\rightarrow\{0,1\}^{k}$ a circuit. For $r\in\{0,1,\cdots,k^{\prime}\}$, we say that $C$ is $r$-sequentially invertible if there is an $r$-invertible circuit sequence that samples from the output distribution of $C$ [^7] The full definition of $\left(r,t,\ell,\varepsilon\right)$-sequential invertibility, which takes into consideration more parameters and not only $r$, is given in Section 4.3. For the needs for this overview, we give the minimal definitions.. An $r$-invertible circuit sequence on $k$ bits is a sequence of pairs of circuits $\left(C_{i,\rightarrow},C_{i,\leftarrow}\right)_{i\in[\ell]}$ such that for every $i\in[\ell]$,
@@
Similarly to the case of the regular Statistical Difference problem $\mathsf{SD}_{a(n),b(n)}$, we have the following promise on the gap of the total variation distance between the circuits. Formally, for $b\in\{0,1\}$ denote
@@
where the randomness sequence is sampled in both cases uniformly $z_{1}\leftarrow\{0,1\}^{r^{b}_{1}(n)},\cdots,z_{\ell}\leftarrow\{0,1\}^{r^{b}_{\ell}(n)}$ for the suitable $b\in\{0,1\}$. We denote the total variation distance between the distributions $D_{0}$, $D_{1}$ with $d$, and it is promised that either $d\leq a(n)$, in which case we say that the input belongs to $\prod_{\text{YES}}$, or we have $d>b(n)$, in which case we determine that the input belongs to $\prod_{\text{NO}}$. The goal of the problem is to decide whether the input belongs to $\prod_{\text{YES}}$ or to $\prod_{\text{NO}}$. As our previous convention for the $\mathsf{SD}$ problem, we also drop the subscript $a(n),b(n)$ to implicitly be any pair of functions with difference $b(n)^{2}-2a(n)+a(n)^{2}$ lower-bounded inverse polynomially, and write $\mathsf{SISD}_{\mathsf{poly},\,r(n)}$.
@@ 2.4 Reducing $\mathsf{GI}$ and $\mathsf{GapCVP}$ to $\mathsf{SISD}_{\mathsf{poly},O(\log(n))}$
In the last part of the Technical Overview we will show that two problems that are believed to be hard for quantum computers, are solvable by a quantum computer with computable order interference, by reducing to the $\mathsf{SISD}_{\mathsf{poly},O(\log(n))}$ problem. In the first part of this section we show the reduction from $\mathsf{GI}$ which is more immediate, and in the second part we show the more involved reduction from $\mathsf{GapCVP}_{O\left(n\sqrt{n}\right)}$. To this end, we will first review the existing reductions from each of the problems to the more general $\mathsf{SD}_{\mathsf{poly}}$ problem (which we don’t know how to solve), and then move to our new reductions, to the easier problem of $\mathsf{SISD}_{\mathsf{poly},O(\log(n))}$.
@@
  • $d\in\mathbb{N}$ is a natural number.
@@
The input comes with a promise: It is either the case that the vector $\mathbf{t}$ is $d$-close to the lattice i.e., $\text{dist}\left(\mathcal{L}_{\mathbf{B}},\mathbf{t}\right)\leq d$, in which case we denote $\left(\mathbf{B},\mathbf{t},d\right)\in\prod_{\text{YES}}$, or the vector is farther than $d\cdot g(n)$ from the lattice i.e., $\text{dist}\left(\mathcal{L}_{\mathbf{B}},\mathbf{t}\right)>d\cdot g(n)$, in which case we denote $\left(\mathbf{B},\mathbf{t},d\right)\in\prod_{\text{NO}}$. In simple words, a solving algorithm for $\mathsf{GapCVP}_{g(n)}$ gets the promise that the vector $\mathbf{t}$ is either close or far away from the lattice $\mathcal{L}_{\mathbf{B}}$, and needs to decide which is the case. The bigger the gap $g(n)$ is, the easier the problem becomes [^9] In particular, it is known that for an exponential $g(n)\approx 2^{n}$ the $\mathsf{GapCVP}_{2^{n}}$ problem is solvable in classical polynomial time [AKS01].. What we will show is that for $g(n)=O\left(n\sqrt{n}\right)$ there is a reduction $\mathsf{GapCVP}_{g(n)}\leq_{p}\mathsf{SISD}_{\mathsf{poly},O(\log(n))}$. The inspiration and starting point of our reduction is the reduction $\mathsf{GapCVP}_{\sqrt{\frac{n}{O\left(\log\left(n\right)\right)}}}\leq_{p}\mathsf{SD}_{\mathsf{poly}}$ by Goldreich and Goldwasser [GG98], which we recall next.
@@
  • Density: We imagine the support of $P$ as an $n$-dimensional shape, by looking at the elements that are farthest away from the origin and thinking on them as a boundary for the shape. The distribution $P$ should be as dense as possible in its shape, which effectively means having a uniform distribution within the shape’s boundaries. The density requirement emerges from the case where $\left(\mathbf{B},\mathbf{t},d\right)\in\prod_{\text{YES}}$ – in that case we get a uniform sample from the shape around the lattice vectors, and the more the shape is sparse, the less our ability to give a good upper bound on the statistical distance between the output distributions of the circuits $C_{0}$, $C_{1}$.
@@
The distribution $P$ we choose is the $n$-dimensional discrete truncated Gaussian distribution $D^{n}_{B}$ (as per Definition 3.12, $B\in\mathbb{N}$ is the truncation parameter, assume it to be $\approx\frac{d\cdot g(n)}{2\cdot\sqrt{n}}$). It is a known fact that $D^{n}_{B}$ is invariant under rotations in $n$-space, which exactly means it does not prefer any direction of the $n$ directions in $\mathbb{Z}^{n}$. In terms of density, the Gaussian distribution is obviously less dense in its shape than the uniform $n$-ball, which is (part of) the reason that we need a bigger gap $g(n)=O\left(n\sqrt{n}\right)$ for which we know how to reduce $\mathsf{GapCVP}_{g(n)}$ to $\mathsf{SISD}$, compared to the better gap $\sqrt{\frac{n}{O\left(\log\left(n\right)\right)}}$ from the previous reduction. As part of the full proof of the reduction (given in Section 7) we analyze the Gaussian distribution’s sparsity in this manner, and the loss of hardness (which translates to the larger $g(n)\approx n\sqrt{n}>>\sqrt{\frac{n}{O\left(\log\left(n\right)\right)}}$) that the sparsity incurs.
@@
Sampling from $D_{B}$ has a plethora of existing techniques and algorithms (e.g., [GPV08, Pei10, MP12, Ros19]), but it is unclear how to show that these can be executed by a sequentially invertible algorithm. In fact, we did not find any sequentially invertible algorithm that samples from a distribution with negligible statistical distance to $D_{B}$. Once we relax our needs to sample from the discrete Gaussian distribution only approximately (that is, with an inverse-polynomial statistical distance), we find a solution. We give the intuitive steps behind our solution in the following.
  • We recall one of the basic theorems of probability theory - The Central Limit Theorem (CLT). The CLT implies that for a distribution $X$ with expectation 00 and standard deviation $\sigma\in\mathbb{R}_{\geq 0}$, when we take the sum of $\kappa$ i.i.d. samples from $X$, that is, $S:=X_{1}+\cdots+X_{\kappa}$, then the normalized sum $\frac{S}{\sqrt{\kappa}}$ approaches the Gaussian distribution (with standard deviation $\sigma$) as $\kappa$ approaches infinity. This helps us because now, instead of sampling from the Gaussian distribution (which we do not know how to do with a sequentially invertible algorithm), we can repeatedly sample from any distribution $X$, which we aim to be sequentially invertible. There are still two inherent barriers for the above approach:
    1. 1.
      While the CLT asserts closeness to the Gaussian distribution for some number $\kappa$ of summands, it does not guarantee a specific rate of convergence, as the rate depends on the chosen distribution $X$. The first barrier implies the efficiency of the reduction.
    2. 2.
      The second barrier is more delicate: The CLT uses normalization (just like the vector normalization used for sampling from the $n$-ball). Since we are dealing with discrete distributions, normalizing means executing integer (that is, rounded) division by $\sqrt{\kappa}$, which is again a non-invertible operation. Without the normalization in the CLT, it is known that it does not always work.
    To overcome both barriers, we need to choose a distribution $X$ with two properties: (1) It can be sampled by a sequentially invertible algorithm, and (2) The rate of convergence of the sum $S$ (and not only the normalized sum $\frac{S}{\sqrt{\kappa}}$) to the Gaussian distribution, is polynomially fast.
  • We observe that the uniform distribution has two properties that work in our favour, for the distribution $X$. It can be shown that if $X$ has a certain discrete continuity property (which the discrete uniform distribution does have), then the un-normalized sum $S$ gives a good convergence rate to the Gaussian distribution. Formally, in Section 7 we show that summing $\kappa$ i.i.d. samples of the uniform distribution over $\{-2^{\beta},\cdots,-1,0,1,\cdots,2^{\beta}\}$ for $2^{\beta}\approx\frac{B}{\sqrt{\kappa}}$ has statistical distance bounded by $O\left(\frac{1}{\sqrt{\kappa}}\right)$ to the discrete Gaussian distribution $D_{B}$.
@@
To conclude, the circuit sequence $C^{0}$ samples a uniformly random lattice vector $\mathbf{v}$, and then adds an approximate discrete Gaussian vector $\mathbf{e}$, where the approximation is done by summing i.i.d. samples from the uniform distribution. The circuit sequence $C^{1}$ is the same, only that we add $\mathbf{t}$ in the end. Both circuit sequences are shown to be $1$-sequentially invertible. We refer the reader to Section 7 for the full details of the proof, including the analysis of the statistical distance between the output distributions of the circuits.